# On the Expressiveness of Visual Prompt Experts

### v5

@@ E.7 Computational Cost
E.8 Interpretive Visualizations
To facilitate a deeper understanding of our method, we provide visualizations to illustrate the advantages of VAPT. Specifically, we use GradCAM [64] to generate attention maps by computing the gradients of a target concept with respect to the model’s final layer. Figure 8 presents examples from five VTAB-1K datasets, namely Sun397, SVHN, Resisc45, Clevr/count, and KITTI/distance, comparing heatmaps generated by VAPT and VPT. These visualizations enable us to examine the regions of the input data to which each technique directs the model’s focus. We observe that VAPT can localize relevant image regions more accurately than VPT. For instance, in the Sun397 dataset, while VPT struggles to capture the complete structure of an object, VAPT succeeds in identifying and highlighting its key features. This enhanced localization indicates VAPT’s stronger ability to capture salient visual patterns. Consequently, VAPT not only improves performance over VPT but also enhances the interpretability of the model by providing more coherent and precise visual explanations.

### v6

@@ E.7 Computational Cost
E.8 Adversarial Robustness
Figure 8: Comparison of VAPT and VPT performance on VTAB-1K CIFAR-100 under PGD adversarial attack.
To evaluate the adversarial robustness of VAPT, we apply the Projected Gradient Descent (PGD) attack (Madry et al., 2017) to VTAB-1K CIFAR-100 test samples. As illustrated in Figure 8, model performance degrades significantly under attack, which is expected given the limited training data and the absence of adversarial-specific training methods (e.g., adversarial fine-tuning). Despite this, VAPT consistently outperforms VPT under adversarial conditions, suggesting that it may confer some robustness benefits. However, these findings are preliminary, and further experiments are required to substantiate the observed trends.
E.9 Interpretive Visualizations
To facilitate a deeper understanding of our method, we provide visualizations to illustrate the advantages of VAPT. Specifically, we use GradCAM (Selvaraju et al., 2017) to generate attention maps by computing the gradients of a target concept with respect to the model’s final layer. Figure 9 presents examples from five VTAB-1K datasets, namely Sun397, SVHN, Resisc45, Clevr/count, and KITTI/distance, comparing heatmaps generated by VAPT and VPT. These visualizations enable us to examine the regions of the input data to which each technique directs the model’s focus. We observe that VAPT can localize relevant image regions more accurately than VPT. For instance, in the Sun397 dataset, while VPT struggles to capture the complete structure of an object, VAPT succeeds in identifying and highlighting its key features. This enhanced localization indicates VAPT’s stronger ability to capture salient visual patterns. Consequently, VAPT not only improves performance over VPT but also enhances the interpretability of the model by providing more coherent and precise visual explanations.
@@ E.9 Interpretive Visualizations
E.10 Multi-Modal Experiments
To investigate the generalizability of Visual Adaptive Prompt Tuning beyond uni-modal visual recognition, we extend our evaluation to a multi-modal setting, specifically image-caption retrieval. This experiment assesses whether the enhanced functional expressiveness of adaptive prompt experts can facilitate better alignment between visual representations and textual semantics within a contrastive learning framework.
Table 21: Image-Caption Retrieval results on ImageNet. We compare VAPT against VPT using a CLIP ViT-B/32 backbone in a 16-shot setting. The models are evaluated using Recall@100 on the validation set. Bold highlights the best results.
Method Recall@100 (%)
Image $\to$ Text Text $\to$ Image Average
VPT 79.05 74.80 76.93
VAPT (Ours) 82.66 77.72 80.19
Improvement +3.61 +2.92 +3.26
[…]
Results. The empirical results, summarized in Table 21, demonstrate that VAPT significantly outperforms VPT in the few-shot cross-modal retrieval task. VAPT achieves an average R@100 of 80.19%, surpassing the VPT baseline by 3.26%. This performance gain is consistent across both retrieval directions, with improvements of 3.61% for Image-to-Text and 2.92% for Text-to-Image retrieval. These findings suggest that the input-conditional nature of VAPT prompt experts allows the visual encoder to dynamically emphasize features that correlate more effectively with natural language descriptions. Consequently, VAPT achieves superior alignment in the multi-modal embedding space compared to static prompting strategies, without necessitating extensive parameter updates or modification of the text encoder.
Appendix F Use of Large Language Models
Large language models were employed solely for editorial purposes, including grammar correction and spelling refinement. They were not used for content generation, data analysis, or the design of experiments.